On October ​1, 2018​, Allan​, Bob​, and Carl formed the​ A, B and C partnership. Allan contributed $ 28800​; Bob​, $ 36000; and Carl​, $ 55200. Allan will manage the​ store; Bob will work in the store​ three-quarters of the​ time; and Carl will not work in the business
Net income for the year ended September 30​, 2019​, is $ 102000. The first $ 40000 is allocated on the basis of relative partner capital balances. The next $ 23000 is based on​ services, with $ 13000 going to Allan and $ 10000 going to Bob. Any remainder is shared equally. ​(Complete all answer boxes. For amounts that are​ $0, make sure to enter​ "0" in the appropriate​ column.)

A, B, and C Allocation of Profits and Losses Allan Bob Carl Total 102000 Net income (loss) Capital allocation: b. Allan Bob Carl Total capital allocation Net income (loss) remaining for allocation Sharing of next portion based on service: Net income (loss) remaining for allocation Remainder shared equally: Net income (loss) remaining for allocation Net income (loss) allocated to each partner ーー ーーーーーーーー ー!ー ーーー ーーーーーーーー
Answer #1
Allan Bob Carl Total
b) Net income 102000
Capital allocation:
Allan = 40000*28800/(28800+36000+55200)= 9600
Allan = 40000*36000/(28800+36000+55200)= 12000
Allan = 40000*55200/(28800+36000+55200)= 18400
Total capital allocation 40000
Net income remaining for allocation 62000
Share of next portion based on service 13000 10000 23000
Net income remaining for allocation 39000
Remainder shared equally 13000 13000 13000 39000
Net income remaining for allocation 0
Net income allocated to each partner 35600 35000 31400 102000